OpenStudy (insa):

in 5 years time a father will be three times as old as his son. four years ago the father was 6 times as old as his son. find their present ages.

random231 (random231):

let the present age of son be x and father be y. y+5=3(x+5) and y-4=6(x-4) solve the equations to get x and y.

OpenStudy (insa):

thx ;)